# Sociodynamics-inspired Adaptive Coalition and Client Selection in Federated Learning

### v1

@@ Appendix B Implementation Details and Further Experiments
In this appendix, we elaborate on the experimental setup, hyperparameters of FedCVR-Bolt , and baselines, including datasets and models. Moreover, Section B.2 addresses the practical implementation of FedCVR-Bolt in the context of increasing model dimensions. Section B.3 provides a detailed analysis of the synthetic dataset structure that simulates synthetic heterogeneous linear regression for FL. Lastly, Section B.4 presents additional results, along with experiments incorporating error bars.
B.1 Datasets and Implementation Details
@@
We evaluate FedCVR-Bolt on various datasets. For a more straightforward scenario, we consider a one-dimensional linear regression under both IID and non-IID distributions, with and without intercept, i.e., $D=1$ and $D=2$ respectively. The dataset construction process is outlined in Section B.3. Subsequently, we evaluate the performance of FedCVR-Bolt using several classification benchmark datasets that are widely used in FL [3, 21], particularly utilizing the Synthetic dataset [3, 12], Federated MNIST [52], and CIFAR-10 [53]. To simulate a heterogeneous environment, we employ the Synthetic(1,1) setting, while the other datasets are partitioned using a Dirichlet distribution with $\alpha=0.1$, by implementing a sampler according to [21]. For the non-IID Synthetic(1,1) scenario, we configured the model to classify 10 distinct classes with an input dimensionality of 60, employing a logistic regression model as specified in [3]. Regarding the MNIST dataset, a multilayer perceptron was utilized, comprising two layers with 200 hidden neurons and ReLU activations. In contrast, the CIFAR-10 experiments involved a convolutional neural network, which consisted of two convolutional and max-pooling layers, followed by two fully connected layers, for generating class predictions.
@@
In all experiments, the dataset is partitioned across $K=100$ clients, followed by dividing the datasets into training and testing sets to assess model performance. The models have been trained with Stochastic Gradient Descent with $S=10$ local epochs, a batch size of $100$, and a learning rate of $\eta=0.01$. We set the number of communication rounds to $T=200$ for the MNIST dataset experiments and $T=100$ for the other datasets. For each round, $P=10$ clients are sampled to participate in the training. The experiments displayed in Tables 1 and 2 utilize a single seed $42$. In contrast, those shown in Tables 3 and 4 employ three different seeds, with $2\sigma$ error bars included to account for variability.
@@
We employed $\gamma_{t}=1/t$ in the training of FedCVR-Bolt to ensure the convergence of Robbins-Monro estimators applied to the covariance [51], followed by the temperature parameter $\beta=1$. In the case of Power-of-choice [12], we obtained $d=2P$ samples and subsequently selected the $P$ that exhibited the highest test loss on the global model. For Active FL [14], we set $\alpha_{1}=0.8$ (which is equivalent to $d=2P$), and similarly utilized the temperature parameter $\alpha_{2}=1$ as our temperature parameter $\beta$, $\alpha_{3}=0$ to achieve comparability with Power-of-choice.
@@
B.2 Practical Implementation of FedCVR-Bolt
Practically, FedCVR-Bolt clusters client according to the observed model $\theta_{k}(t)$. In order to induce a minor drift and facilitate the initiation of the dynamic clustering process at each iteration, we implement a uniformly random client selection strategy during the initial 30 rounds, enabling the observation of local adaptations of the global model across various sampled clients. The selection of 30 rounds offers a balanced approach between exploring client heterogeneity and subsequently leveraging the clustering structure.
Similar to numerous FL methods based on local updates – such as [9] – executing operations on the full model becomes increasingly burdensome as the model size increases. Furthermore, from a statistical standpoint, the early feature extraction layers (such as convolutional or embedding layers) frequently contribute less discriminative information for clustering and may introduce superfluous noise along with computational overhead. Hence, when managing larger models, we confine the covariance computation to the parameters of the final fully connected layer. If this layer remains excessively large, we further diminish dimensionality by randomly sampling a subset of its weights. Specifically, for the MNIST dataset, we utilize the entire final layer; whereas for CIFAR-10, where the final layer has higher dimensionality, we sample $D=300$ weights. This approach achieves an advantageous compromise between computational efficiency and representational adequacy.
B.3 Synthetic Linear Regression
@@
B.4 Further Experiments
In this section, we present $2\sigma$ error bars that were obtained by executing the algorithms with three distinct random seeds, referring to Tables 4 and 2. In Table 3, it is evident that the results pertaining to synthetic regression exhibit greater variability, attributable to the slight alterations in the data distribution influenced by the seed, thereby engendering a constant error across different seeds. Nonetheless, FedCVR-Bolt demonstrates superior performance, including in the IID scenario. With regard to the classification results illustrated in Table 4, FedCVR-Bolt consistently surpasses the baselines, while also exhibiting reduced errors, thus indicating that coalition formation remains robust against statistical perturbations.
[…]
Regression Model Heterogeneity FedAvg Power-of-Choice AFL FedCVR-Bolt $y=\theta_{1}x$ IID 6.0171 $\pm$ 2.4884 6.3587 $\pm$ 2.0280 6.1656 $\pm$ 2.5821 6.0136$\pm$ 2.5170 Non-IID 57.1336 $\pm$ 24.1335 56.9701 $\pm$ 24.7835 56.8328 $\pm$ 25.7835 54.7102 $\pm$ 24.8973 $y=\theta_{0}+\theta_{1}x$ IID 0.2142$\pm$0.3520 0.2143 $\pm$ 0.3533 0.2134 $\pm$ 0.3503 0.2085 $\pm$ 0.3386 Non-IID 1.5875 $\pm$ 0.7291 1.6742 $\pm$ 0.6820 1.7023 $\pm$ 0.7642 1.5127 $\pm$ 0.7336
[…]
Dataset FedAvg Power-of-Choice Active FL FedCVR-Bolt
Synthetic (1,1) 72.67 $\pm$ 1.99 78.13 $\pm$ 2.59 78.06 $\pm$ 2.60 79.35 $\pm$ 1.16
MNIST 86.93 $\pm$ 1.02 84.86 $\pm$ 5.95 90.07 ±$\pm 1.06$ 90.82 $\pm$ 0.85
CIFAR-10 52.84 $\pm$ 1.71 49.92 $\pm$ 0.85 51.13 $\pm$ 6.67 56.56 $\pm$ 0.76
Appendix C Privacy of FedCVR-Bolt
The privacy of client data in FedCVR-Bolt algorithm is preserved by adhering to the fundamental principles of FL: raw data remains on client devices and is never transmitted to the server. Clients selected from the set $\mathcal{P}_{t}$ perform local updates using the global model $\theta_{gl}(t)$ to produce their updated local models $\theta_{j_{p}}(t+1)$. Moreover, clients subsequently communicate to the server only these model parameters, not the underlying private data. To further protect such parameters during aggregation steps (e.g., for the computation of $\theta_{gl}(t+1)$), it is possible to employ secure aggregation protocols [57]. While the distinctive server-side operations of FedCVR-Bolt , such as clustering based on individual models $\{\theta_{k}(t)\}$ and deriving $\bar{\theta}_{k}(t+1)$ from specific client model updates $\theta_{j_{p}}(t+1)$, require the server to access these individual parameters for its core functionality, this architectural choice is consistent with federated architectures where the central server orchestrates the learning process using the model parameters received from the clients [58] for advanced tasks like personalization or clustering [7]. Thus, FedCVR-Bolt ensures foundational data privacy by localizing data processing, and the principles of secure aggregation offer a complementary mechanism for enhancing model confidentiality during aggregation where the algorithm’s direct need for individual parameter access permits.

### v2

@@ Appendix C Implementation Details and Further Experiments
In this appendix, we focus on the experimental setup, hyperparameters of FedCVR-Bolt , and baselines, including datasets and models. Moreover, Section C.2 addresses the practical implementation of FedCVR-Bolt in the context of increasing model dimensions. Finally, Section C.3 provides a detailed analysis of the synthetic dataset structure that simulates synthetic heterogeneous linear regression for FL.
C.1 Datasets and Implementation Details
@@
We evaluate FedCVR-Bolt on various datasets. For a more straightforward scenario, we consider a one-dimensional linear regression under both IID and non-IID distributions, with and without intercept, i.e., $D=1$ and $D=2$ respectively. The dataset construction process is outlined in Section C.3. Subsequently, we evaluate the performance of FedCVR-Bolt using several classification benchmark datasets that are widely used in FL (Li et al., 2020; Caldas et al., 2018), particularly utilizing the Synthetic dataset (Li et al., 2020; Cho et al., 2022), Federated MNIST (LeCun, 1998), and CIFAR-10 (Krizhevsky, 2009). To simulate a heterogeneous environment, we employ the Synthetic(1,1) setting, while the other datasets are partitioned using a Dirichlet distribution with $\alpha=0.1$, by implementing a sampler according to Caldas et al. (2018). Regarding the MNIST dataset, a multilayer perceptron was utilized, comprising two layers with 200 hidden neurons and ReLU activations. In contrast, the CIFAR-10 and CIFAR-100 experiments involved a convolutional neural network, which consisted of two convolutional and max-pooling layers, followed by two fully connected layers, for generating class predictions.
@@
In all experiments, the dataset is partitioned across $K=100$ clients, followed by dividing the datasets into training and testing sets to assess model performance. The models have been trained with Stochastic Gradient Descent with $S=10$ local epochs, a batch size of $100$, and a learning rate of $\eta=0.01$. We set the number of communication rounds to $T=200$ for the MNIST dataset experiments and $T=100$ for the other datasets. For each round, $P=10$ clients are sampled to participate in the training.
@@
We employed $\gamma_{t}=1/t$ in the training of FedCVR-Bolt to ensure the convergence of Robbins-Monro estimators applied to the covariance (Robbins & Monro, 1951), followed by the temperature parameter $\beta=1$. In the case of Power-of-choice (Cho et al., 2022), we obtained $d=2P$ samples and subsequently selected the $P$ that exhibited the highest test loss on the global model. For Active FL (Goetz et al., 2019), we set $\alpha_{1}=0.8$ (which is equivalent to $d=2P$), and similarly utilized the temperature parameter $\alpha_{2}=1$ as our temperature parameter $\beta$, $\alpha_{3}=0$ to achieve comparability with Power-of-choice.
@@
C.2 Practical Implementation of FedCVR-Bolt
Practically, FedCVR-Bolt clusters client according to the observed model $\theta_{k}(t)$. To further stabilize the initial estimation phase, we implement a uniformly random sampling strategy during the initial 30 rounds, enabling the observation of local adaptations of the global model across various sampled clients. The selection of 30 rounds offers a balanced approach between exploring client heterogeneity and subsequently leveraging the clustering structure. This initial phase is dedicated to observing a diverse set of client behaviors, which allows the Robbins-Monro estimator to build a more reliable and stable baseline covariance structure before our variance-reduction sampling policy becomes active (Grimmett & Stirzaker, 2001).
Similar to numerous FL methods based on local updates — such as Ghosh et al. (2020) — executing operations on the full model becomes increasingly burdensome as the model size increases. Furthermore, from a statistical standpoint, the early feature extraction layers (such as convolutional or embedding layers) frequently contribute less discriminative information for clustering and may introduce superfluous noise along with computational overhead. Hence, when managing larger models, we confine the covariance computation to the parameters of the final fully connected layer. If this layer remains excessively large, we further diminish dimensionality by randomly sampling a subset of its weights. Specifically, for the MNIST dataset, we utilize the entire final layer; whereas for CIFAR-10, where the final layer has higher dimensionality, we sample $D=300$ weights. This approach achieves an advantageous compromise between computational efficiency and representational adequacy.
C.3 Synthetic Linear Regression
@@
Appendix D Privacy of FedCVR-Bolt
The privacy of client data in FedCVR-Bolt algorithm is preserved by adhering to the fundamental principles of FL: raw data remains on client devices and is never transmitted to the server. Clients selected from the set $\mathcal{P}_{t}$ perform local updates using the global model $\theta_{gl}(t)$ to produce their updated local models $\theta_{j_{p}}(t+1)$. Moreover, clients subsequently communicate to the server only these model parameters, not the underlying private data. To further protect such parameters during aggregation steps (e.g., for the computation of $\theta_{gl}(t+1)$), it is possible to employ secure aggregation protocols (Bonawitz et al., 2016). While the distinctive server-side operations of FedCVR-Bolt , such as clustering based on individual models $\{\theta_{k}(t)\}$ and deriving $\bar{\theta}_{k}(t+1)$ from specific client model updates $\theta_{j_{p}}(t+1)$, require the server to access these individual parameters for its core functionality, this architectural choice is consistent with federated architectures where the central server orchestrates the learning process using the model parameters received from the clients (Kairouz et al., 2021) for advanced tasks like personalization or clustering (Smith et al., 2017). Thus, FedCVR-Bolt ensures foundational data privacy by localizing data processing, and the principles of secure aggregation offer a complementary mechanism for enhancing model confidentiality during aggregation where the algorithm’s direct need for individual parameter access permits.